question:
In a triangle the measure of the second angle is one-half the measure of the first angle. The third angle's measure if 20 degrees less than the measure of the first angle. What are the masures of the angles?
Please help! daddy's girl

Answer by askmemath(368)

Let the first angle = X
Therefore the 2nd angle becomes X/2 = 0.5X
And the 3rd angle becomes X-20
We know that sum of angles in a triangle = 180degrees
That means
X + 0.5X + X-20 = 180
Adding 20 on both sides
2.5X = 200
Dividing by 2.5 on both sids
X = 80
which means 2nd angle = 80/2 = 40
and the third angle = 80-20 = 60
Check:
Sum of the 3 angles
80+40+60 = 180degrees
Yay!!!